# Boost Post-Training Quantization via Null Space Optimization for Large Language Models

### v2

@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023) have demonstrated remarkable performance across various tasks in recent years. However, billions of parameters in LLMs incurs significant storage and inference overheads. To address these issues, quantization (Li et al., 2021; Dettmers et al., 2022; Nagel et al., 2020), which reduces memory requirements and accelerates inference by converting high-precision values in LLMs into low-bit representations, receives tremendous attention. Existing quantization approaches can be divided into Quantization-aware Training (QAT) (Liu et al., 2023; Xu et al., 2024; Du et al., 2024) and Post-training Quantization (PTQ) (Xiao et al., 2023; Lin et al., 2024; Yuan et al., 2023) according to the pipeline. Between them, PTQ is more popular in LLM community because of its efficiency and resource-friendliness. Currently, PTQ techniques for LLMs can achieve lossless performance at 4-bit (Frantar et al., 2022; Shao et al., 2023; Zhao et al., 2024) and high-accuracy inference at 1.61-3 bits (Chee et al., 2023; Zhao et al., 2025b; Huang et al., 2024).
@@
Despite the development of various advanced quantization strategies, they all suffer from a shared limitation. As known, all existing methods are uniformly motivated by the goal of minimizing quantization error. For example, the weight-only quantization error is formulated by $\|WX-W_{q}X\|_{2}^{2}$ (Zhao et al., 2025a). From a numerical perspective, regardless of how effective the quantization method is, the error is fundamentally unavoidable. Without additional constraints, such numerical error ($W-W_{q}$) is bound to negatively affect the final output. Recognizing this limitation, we systematically consider the following question: Given the inevitability of numerical error, is there a way to alleviate its impact on the final quantization error?
@@
Aiming at this question, in this paper we propose a pioneering perspective that quantization error can be effectively alleviated through the theoretical properties of Null Space (Coleman and Pothen, 1986, 1987; Ravfogel et al., 2020). Specifically, for weight-only quantization, we aim to prove that as long as the weight perturbation ($W-W_{q}$) lies in the null space of the input activations, the final quantization error will be alleviated effectively ($\|(W-W_{q})X\|_{2}^{2}\approx 0$). Here we select a straightforward method to prove this perspective: constructing an approximate projection for quantized model that maps the numerical error into the null space. If performance enhancements appear, this insight will be proved to be effective.
@@
During validation, we face two main challenges. Typically, computing the null space projection relies on SVD decomposition (Wang et al., 2024), where the left singular vectors corresponding to zero singular values span the null space (Fang et al., 2024; Tang et al., 2025). However, the large matrix dimensions make SVD computation extremely expensive, and the singular values of activation matrices rarely exhibit exact zeros. To address these issues, we propose an efficient and accurate approximation method to get the null space projection $\Delta$ based on the Prefix-Suffix Sum Ratio of singular values. In addition, since there only exists $W_{q}$ during inference for a quantized model, simply applying $(W-W_{q})\Delta X\approx 0$ is not practically meaningful, and storing $\Delta$ would incur additional memory overhead. Considering this, we reformulate the null space optimization as solving $W-\alpha W_{q}=(W-W_{q})\Delta$ and derive a closed-form solution for the equivalent projection vector $\alpha$, where $\alpha$ can be easily absorbed into the scaling factors to avoid extra memory costs while achieving null space optimization.
@@ 2.1 Quantization for LLMs
Quantization is one of the most widely studied model compression techniques (Frantar and Alistarh, 2023; Gou et al., 2021; Hu et al., 2022), offering both high performance and high compression ratios. According to the pipeline, existing approaches fall into two categories: Quantization-aware training (QAT) and Post-training quantization (PTQ). QAT (Liu et al., 2023; Wang et al., 2023; Ma et al., 2024) requires training from scratch and updating the weights, which can achieve higher performance generally but the increased training overheads significantly hinders its development.
In contrast, PTQ only need a small scale calibration set to get effective quantization parameters without retraining, so its efficiency and resource-friendliness have made it more popular in LLM community. GPTQ Frantar et al. (2022) leverages second-order information to dynamically update the remaining weights during quantization. AWQ (Lin et al., 2024) assesses the saliency of different weight channels based on the input activations and allocates appropriate scaling factors accordingly. QuIP (Chee et al., 2023) utilizes incoherence preprocessing to transform the weights, achieving high performance at 2-bit. Despite their success, all existing methods suffer from inevitable performance degradation caused by conventional quantization error formulation, demonstrating imperative extra optimization constraint.
@@ 2.2 Null Space Constraint Learning
Null space is a classical concept in linear algebra which is extensively studied in mathematics (Frittelli et al., 1997), and recently it is increasingly applied in machine learning (Zhang et al., 2016). Adam-NSCL (Wang et al., 2021) introduces null space optimization into continual learning by forcing the network parameter update lying in the null space of the input feature to balance plasticity and stability. LoRA-Null (Tang et al., 2025) builds adapters initialized from the null space of the pretrained knowledge acitivation to encounter catastrophic forgetting in model finetuning. In the area of knowledge editing, AlphaEdit (Fang et al., 2024) leverages this theoretical insights to balance knowledge-update error and knowledge-preservation error. Inspired by these advances, for the first time we introduce null space theory into model quantization to mitigate the impact of quantization error and further enhance quantization performance.
@@ 3.2 Efficient and Accurate Approximation for Null Space Projection
Following existing methods (Wang et al., 2021; Fang et al., 2024), the layer-wise null space of input activation $X$ can be modeled as that of their uncentered covariance matrix $XX^{T}$ to guarantee stability, whose null space is equal to that of $X$ (please refer to Appendix for detailed proof). Subsequently, the first step of the conventional method for conducting null space projection is to apply SVD decomposition to $XX^{T}$:
@@
Although the two are mathematically related, in practice, computing SVD requires multiple Householder transformations and leverages implicit QR iterations to approximate the singular values, which result in a more complex computation pipeline. In contrast, eigenvalue decomposition allows for faster eigenvalue computation via the QR algorithm (Watkins, 1982), and efficient recovery of eigenvectors through inverse iteration. Moreover, PyTorch’s backend acceleration framework (Paszke, 2019) employs a divide-and-conquer strategy specifically optimized for eigenvalue decomposition of real symmetric matrices to achieve higher computational efficiency. Therefore, we suggest estimating the null space for LLMs by eigenvalue decomposition via QR-based iteration instead of conventional SVD, which significantly improves decomposition efficiency. In Table 4 we report the improvements in decomposition efficiency and the corresponding accuracy comparison, demonstrating its superiority.
@@
It is worth noting that (Wang et al., 2021) estimates the null space by selecting singular values that are 50 times larger than the smallest one. While effective for small-scale image features in CNNs, their approach fails to provide reliable null space estimation for LLMs. Detailed comparison among PyTorch, Adam-NSCL and ours are provided in Table 5, demonstrating our method consistently leads to better performance.
@@ 4.1 Experimental Setup
Our experiments are primarily conducted on LLaMA3 (8B, 70B), LLaMA3.1 (8B, 70B) and LLaMA3.3 (70B) (Grattafiori et al., 2024), as they are currently the most popular and widely applied open-sourced LLMs. In addition, three prominent star LLMs (DeepSeekMoE-16B (Dai et al., 2024), Qwen2.5-32B (Yang et al., 2024) and Qwen3-32B (Team, 2025)) are also included for evaluation. We integrate our Q2N with GPTQ (Frantar et al., 2022), the broadest practically deployed LLMs PTQ method, to valid different performance aspects. To ensure generality, we also select four baseline methods (QuIP (Chee et al., 2023), PB-LLM (Shang et al., 2023), LeanQuant (Zhang and Shrivastava, 2024) and QuaRot (Ashkboos et al., 2024)) which represent diverse quantization strategies according to (Zhao et al., 2025a).
@@
According to Section 3.1, quantization error becomes more pronounced at lower bit, so our null space optimization focus on 2(g128)/3-bit scenario. For accurate null space approximation and the closed-form solution of the projection vector, the threshold $t$ is set to $0.1$ and the regularization coefficient $\lambda$ is set to $0.2$ (optimal in most cases, please refer to Appendix C for detailed). During quantization, our calibration set consists of 128 random 2048 token-segments from WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020). All procedures are deployed on 1 NVIDIA A800-80G GPU.
@@
We evaluate language generation capability (perplexity $\downarrow$) and downstream reasoning capability (zero-shot accuracy $\uparrow$) for the optimized quantized LLMs. For language generation tasks, the test data comes from WikiText2, PTB (Marcus et al., 1994) and C4. Downstream reasoning tasks includes ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Race (Lai et al., 2017), MMLU (Hendrycks et al., 2020), PIQA (Bisk et al., 2020) and WinoGrande (Sakaguchi et al., 2021), using the open-sourced toolkit lm-evaluation-harness (Gao et al., 2024) to evaluate.
@@ 4.3 Performance Enhancements on Baseline Methods within Different Strategies (RQ2)
To valid whether quantization strategy will influence the scalability, followed by (Zhao et al., 2025a) we select three SOTA baseline methods from different PTQ strategies to combine with our Q2N: QuIP (Chee et al., 2023) represents rotation-based strategy, PB-LLM (Shang et al., 2023) typifies the combination of salience-based and compensation-based strategy, and LeanQuant (Zhang and Shrivastava, 2024) exemplifies the combination of optimization-based and compensation-based strategy. As indicated by Figure 2, all three baselines experience notable performance improvements. For example, QuIP achieves a $7.3$ reduction in perplexity on WikiText2 at 2-bit, LeanQuant improves inference accuracy on HellaSwag by $10.7\%$.
We also extend the scope to weight-activation (WA) PTQ scenario. Specifically, we select QuaRot (Ashkboos et al., 2024), the state-of-the-art WA method, to perform W4A4 quantization on LLaMA families, and then use Q2N to optimize the quantized weights after its original process. As demonstrated in Table 3, our Q2N further pushes the performance limits at each metric. Therefore, we can answer RQ2 that null space optimization enables performance improvements on diverse quantization strategies, including the SOTA weight-activation method.
@@ 4.5 Accurate Null Space Approximation vs Adam-NSCL & PyTorch (RQ4)
In Section 3.2, we introduce an accurate null space approximation method based on the Prefix-Suffix Sum Ratio of singular values to address the challenge that the activations in LLMs rarely contain exact zero singular values. Previously, Adam-NSCL (Wang et al., 2021) computes the null space in continual learning for CNN-based image classification. In addition, PyTorch also provides a built-in matrix rank estimation function Torch.linalg.matrix_rank(), which implicitly ignores small singular values. In Table 5, we compare the three approaches on LLaMA3.1-8B (3-bit, C4 calibration), where our null space approximation method used in Q2N consistently outperforms the others, with particularly strong gains on HellaSwag, exceeding both alternatives by nearly $10\%$. Notably, all null space based methods improve upon naive GPTQ, indirectly validating the importance of integrating null space optimization with LLMs quantization.
@@ A.2 Proof for Null Space Projection $\Delta=U_{1}U_{1}^{T}$
which indicates that the columns of $U_{1}$ span the null space for $XX^{T}$. According to [Meyer, 2023] (Eq. 5.13.4), the orthogonal projector of $XX^{T}$ can be elaborated as:

### v3

@@ 1 Introduction
Large language models (LLMs) [39, 1, 3] have demonstrated remarkable performance across various tasks in recent years. However, billions of parameters in LLMs incurs significant storage and inference overheads. To address these issues, quantization [23, 10, 30], which reduces memory requirements and accelerates inference by converting high-precision values in LLMs into low-bit representations, receives tremendous attention. Existing quantization approaches can be divided into Quantization-aware Training (QAT) [25, 45, 11] and Post-training Quantization (PTQ) [44, 24, 47] according to the pipeline. Between them, PTQ is more popular in LLM community because of its efficiency and resource-friendliness. Currently, PTQ techniques for LLMs can achieve lossless performance at 4-bit [14, 36, 51] and high-accuracy inference at 1.61-3 bits [5, 53, 21].
@@
Despite the development of various advanced quantization strategies, they all suffer from a shared limitation. As known, all existing methods are uniformly motivated by the goal of minimizing quantization error. For example, the weight-only quantization error is formulated by $\|WX-W_{q}X\|_{2}^{2}$ [52]. From a numerical perspective, regardless of how effective the quantization method is, the error is fundamentally unavoidable. Without additional constraints, such numerical error ($W-W_{q}$) is bound to negatively affect the final output. Recognizing this limitation, we systematically consider the following question: Given the inevitability of numerical error, is there a way to alleviate its impact on the final quantization error?
@@
Aiming at this question, in this paper we propose a pioneering perspective that quantization error can be effectively alleviated through the theoretical properties of Null Space [7, 8, 33]. Specifically, for weight-only quantization, we aim to prove that as long as the weight perturbation ($W-W_{q}$) lies in the null space of the input activations, the final quantization error will be alleviated effectively ($\|(W-W_{q})X\|_{2}^{2}\approx 0$). Here we select a straightforward method to prove this perspective: constructing an approximate projection for quantized model that maps the numerical error into the null space. If performance enhancements appear, this insight will be proved to be effective.
@@
During validation, we face two main challenges. Typically, computing the null space projection relies on SVD decomposition [42], where the left singular vectors corresponding to zero singular values span the null space [12, 37]. However, the large matrix dimensions make SVD computation extremely expensive, and the singular values of activation matrices rarely exhibit exact zeros. To address these issues, we propose an efficient and accurate approximation method to get the null space projection $\Delta$ based on the Prefix-Suffix Sum Ratio of singular values. In addition, since there only exists $W_{q}$ during inference for a quantized model, simply applying $(W-W_{q})\Delta X\approx 0$ is not practically meaningful, and storing $\Delta$ would incur additional memory overhead. Considering this, we reformulate the null space optimization as solving $W-\alpha W_{q}=(W-W_{q})\Delta$ and derive a closed-form solution for the equivalent projection vector $\alpha$, where $\alpha$ can be easily absorbed into the scaling factors to avoid extra memory costs while achieving null space optimization.
@@ 2.1 Quantization for LLMs
Quantization is one of the most widely studied model compression techniques [13, 17, 20], offering both high performance and high compression ratios. According to the pipeline, existing approaches fall into two categories: Quantization-aware training (QAT) and Post-training quantization (PTQ). QAT [25, 40, 26] requires training from scratch and updating the weights, which can achieve higher performance generally but the increased training overheads significantly hinders its development.
In contrast, PTQ only need a small scale calibration set to get effective quantization parameters without retraining, so its efficiency and resource-friendliness have made it more popular in LLM community. GPTQ [14] leverages second-order information to dynamically update the remaining weights during quantization. AWQ [24] assesses the saliency of different weight channels based on the input activations and allocates appropriate scaling factors accordingly. QuIP [5] utilizes incoherence preprocessing to transform the weights, achieving high performance at 2-bit. Despite their success, all existing methods suffer from inevitable performance degradation caused by conventional quantization error formulation, demonstrating imperative extra optimization constraint.
@@ 2.2 Null Space Constraint Learning
Null space is a classical concept in linear algebra which is extensively studied in mathematics [15], and recently it is increasingly applied in machine learning [49]. Adam-NSCL [41] introduces null space optimization into continual learning by forcing the network parameter update lying in the null space of the input feature to balance plasticity and stability. LoRA-Null [37] builds adapters initialized from the null space of the pretrained knowledge acitivation to encounter catastrophic forgetting in model finetuning. In the area of knowledge editing, AlphaEdit [12] leverages this theoretical insights to balance knowledge-update error and knowledge-preservation error. Inspired by these advances, for the first time we introduce null space theory into model quantization to mitigate the impact of quantization error and further enhance quantization performance.
@@ 3.2 Efficient and Accurate Approximation for Null Space Projection
Following existing methods [41, 12], the layer-wise null space of input activation $X$ can be modeled as that of their uncentered covariance matrix $XX^{T}$ to guarantee stability, whose null space is equal to that of $X$ (please refer to Appendix for detailed proof). Subsequently, the first step of the conventional method for conducting null space projection is to apply SVD decomposition to $XX^{T}$:
@@
Although the two are mathematically related, in practice, computing SVD requires multiple Householder transformations and leverages implicit QR iterations to approximate the singular values, which result in a more complex computation pipeline. In contrast, eigenvalue decomposition allows for faster eigenvalue computation via the QR algorithm [43], and efficient recovery of eigenvectors through inverse iteration. Moreover, PyTorch’s backend acceleration framework [31] employs a divide-and-conquer strategy specifically optimized for eigenvalue decomposition of real symmetric matrices to achieve higher computational efficiency. Therefore, we suggest estimating the null space for LLMs by eigenvalue decomposition via QR-based iteration instead of conventional SVD, which significantly improves decomposition efficiency. In Table 4 we report the improvements in decomposition efficiency and the corresponding accuracy comparison, demonstrating its superiority.
@@
It is worth noting that [41] estimates the null space by selecting singular values that are 50 times larger than the smallest one. While effective for small-scale image features in CNNs, their approach fails to provide reliable null space estimation for LLMs. Detailed comparison among PyTorch, Adam-NSCL and ours are provided in Table 5, demonstrating our method consistently leads to better performance.
@@ 4.1 Experimental Setup
Our experiments are primarily conducted on LLaMA3 (8B, 70B), LLaMA3.1 (8B, 70B) and LLaMA3.3 (70B) [18], as they are currently the most popular and widely applied open-sourced LLMs. In addition, three prominent star LLMs (DeepSeekMoE-16B [9], Qwen2.5-32B [46] and Qwen3-32B [38]) are also included for evaluation. We integrate our Q2N with GPTQ [14], the broadest practically deployed LLMs PTQ method, to valid different performance aspects. To ensure generality, we also select four baseline methods (QuIP [5], PB-LLM [35], LeanQuant [50] and QuaRot [2]) which represent diverse quantization strategies according to [52].
@@
According to Section 3.1, quantization error becomes more pronounced at lower bit, so our null space optimization focus on 2(g128)/3-bit scenario. For accurate null space approximation and the closed-form solution of the projection vector, the threshold $t$ is set to $0.1$ and the regularization coefficient $\lambda$ is set to $0.2$ (optimal in most cases, please refer to Appendix C for detailed). During quantization, our calibration set consists of 128 random 2048 token-segments from WikiText2 [28] and C4 [32]. All procedures are deployed on 1 NVIDIA A800-80G GPU.
@@
We evaluate language generation capability (perplexity $\downarrow$) and downstream reasoning capability (zero-shot accuracy $\uparrow$) for the optimized quantized LLMs. For language generation tasks, the test data comes from WikiText2, PTB [27] and C4. Downstream reasoning tasks includes ARC [6], HellaSwag [48], Race [22], MMLU [19], PIQA [4] and WinoGrande [34], using the open-sourced toolkit lm-evaluation-harness [16] to evaluate.
@@ 4.3 Performance Enhancements on Baseline Methods within Different Strategies (RQ2)
To valid whether quantization strategy will influence the scalability, followed by [52] we select three SOTA baseline methods from different PTQ strategies to combine with our Q2N: QuIP [5] represents rotation-based strategy, PB-LLM [35] typifies the combination of salience-based and compensation-based strategy, and LeanQuant [50] exemplifies the combination of optimization-based and compensation-based strategy. As indicated by Figure 2, all three baselines experience notable performance improvements. For example, QuIP achieves a $7.3$ reduction in perplexity on WikiText2 at 2-bit, LeanQuant improves inference accuracy on HellaSwag by $10.7\%$.
We also extend the scope to weight-activation (WA) PTQ scenario. Specifically, we select QuaRot [2], the state-of-the-art WA method, to perform W4A4 quantization on LLaMA families, and then use Q2N to optimize the quantized weights after its original process. As demonstrated in Table 3, our Q2N further pushes the performance limits at each metric. Therefore, we can answer RQ2 that null space optimization enables performance improvements on diverse quantization strategies, including the SOTA weight-activation method.
@@ 4.5 Accurate Null Space Approximation vs Adam-NSCL & PyTorch (RQ4)
In Section 3.2, we introduce an accurate null space approximation method based on the Prefix-Suffix Sum Ratio of singular values to address the challenge that the activations in LLMs rarely contain exact zero singular values. Previously, Adam-NSCL [41] computes the null space in continual learning for CNN-based image classification. In addition, PyTorch also provides a built-in matrix rank estimation function Torch.linalg.matrix_rank(), which implicitly ignores small singular values. In Table 5, we compare the three approaches on LLaMA3.1-8B (3-bit, C4 calibration), where our null space approximation method used in Q2N consistently outperforms the others, with particularly strong gains on HellaSwag, exceeding both alternatives by nearly $10\%$. Notably, all null space based methods improve upon naive GPTQ, indirectly validating the importance of integrating null space optimization with LLMs quantization.
@@ A.2 Proof for Null Space Projection $\Delta=U_{1}U_{1}^{T}$
which indicates that the columns of $U_{1}$ span the null space for $XX^{T}$. According to [29] (Eq. 5.13.4), the orthogonal projector of $XX^{T}$ can be elaborated as: